# Grasping a Handful: Sequential Multi-Object Dexterous Grasp Generation

### v4

@@ I Introduction
Dexterous grasping has been extensively studied in robotics [1, 2]. The field has evolved through two primary research directions. First, numerous methods have been developed for generating robotic grasps [3, 4, 5, 6, 7, 8], from analytical approaches [7] to data-driven generative models including Generative Adversarial Networks [5], diffusion models [4, 3], and Variational Autoencoders [8, 6]. Second, a complementary line of research has focused on understanding human grasping behaviors, including intention recognition [9], real-time tracking of hand-object interactions [10], grasp taxonomy classification [11], and learning task-oriented grasping strategies from human demonstrations [12].
The primary strategy in dexterous grasp generation has been to bring the robotic hand close to the object and then simultaneously envelop it with all fingers. While this strategy often results in efficient and successful grasp generation, it simplifies dexterous grasping to resemble parallel-jaw grasping, thereby underutilizing the many DoF of multi-fingered robotic hands [7].
@@
We experimentally evaluate SeqGrasp, SeqDiffuser, and the state-of-the-art simultaneous multi-object grasping method MultiGrasp [13] in simulation and on physical hardware. The simulation results revealed that SeqGrasp and SeqDiffuser perform on par with MultiGrasp for picking one or two objects while outperforming it when picking three to four objects. Moreover, SeqDiffuser demonstrates superior efficiency, generating 256 grasps within one second compared to approximately 1000 seconds for SeqGrasp and MultiGrasp. For the real-world experiments, we replicated the grasp sequences proposed by the methods on a real Allegro Hand attached to a Franka Panda Robot. These results align with the simulation findings, demonstrating that SeqGrasp reaches a 43.33% higher grasp success rate than MultiGrasp.
@@
MOG [14] Human&Barrett Sim./Real 11 \ 28K Simu./Seq. \
Grasp’Em [13] Shadow Sim. 8 36 90K Simultaneous 2
@@ II-A Dexterous Grasping
Analytical methods. Early research in dexterous grasping generated stable grasps by optimizing a grasp quality metric such as the force-closure metric [15, 16]. Although these methods are theoretically sound, they are computationally demanding because (i) the many DoF dexterous hands cause high-dimensional search spaces [17], and (ii) the quality metrics are expensive to compute [18]. Consequently, some methods have focused on reducing the search space by imposing constraints on the hand [19, 20] or restricting it to joint configuration subspaces [17]. Another line of work has proposed a computationally cheap differentiable force-closure estimator [18, 7], which has the advantage of being hand-agnostic. This work extends the differentiable force-closure measure from [7] to sequential multi-object grasping.
Data-driven methods. Recent advancements in machine learning have significantly improved dexterous grasp generation [3, 21, 4, 6]. Nowadays, deep generative models can generate thousands of dexterous grasps on previously unseen and partially observed objects within seconds [3, 21, 4, 6], something analytical methods cannot. Still, only a few data-driven dexterous grasping methods have been developed for multi-object grasping [22, 13]. In this work, we train a new diffusion-based sequential multi-object grasp sampler SeqDiffuser inspired from [3] on our own optimization-generated sequential multi-object dataset SeqDataset.
@@ II-B Multi-Object Grasping
Multi-object grasping presents unique challenges due to the complex multi-object interactions and the high-dimensional configuration space spanned by the hand and the objects. Some prior parallel-jaw multi-object grasping methods [23, 24] explored multi-object push grasps where scattered objects are first pushed together to facilitate multi-object grasping. However, these methods are limited by their reliance on simple shape primitives and parallel-jaw grippers. In comparison, our work can handle objects of diverse shapes and sizes.
A few works have addressed dexterous multi-object grasping [13, 22, 25] where [13] targets simultaneous multi-object grasping while [22, 25] targets sequential multi-object grasping. Li et al. [13] proposed MultiGrasp a two-stage simultaneous multi-object dexterous grasping framework where a generative grasp sampler proposed poses to simultaneously pick many objects, followed by a learned policy for executing the pick. Simultaneously grasping multiple objects is an efficient strategy because it avoids repeated hand repositioning, which reduces total execution time, but the main limitation is that objects must be spatially close and of similar size and shape. In comparison, our method can handle scattered objects of different shapes and sizes by sequentially picking one at a time. The other works that do sequential multi-object grasping [22, 25] restrict the grasping to a maximum of two objects [22] or to primitive object shapes such as cylinders or spheres [25]. In comparison, our method can handle up to four objects of complex shapes and sizes. We also use our method to collect the largest sequential multi-object grasping dataset to date.
@@ II-C Datasets for Dexterous Grasping
Large-scale dexterous grasp datasets [5, 7, 26, 14, 13, 21, 27, 28] have significantly advanced the training of data-driven dexterous grasping methods. However, most of these datasets target single-object grasping [5, 7, 26, 21, 27, 28], with only a few for multi-object grasping [14, 13]. As shown in Table I, these existing multi-object datasets are small and predominantly focus on simultaneous rather than sequential grasping. Therefore, we collect the new large-scale sequential multi-object grasping dataset SeqDataset [^1] The dataset is available at https://yulihn.github.io/SeqGrasp/. using our method SeqGrasp. SeqDataset is, to date, by far the largest multi-object grasping dataset.
@@ III Problem Formulation
To contrast, simultaneous multi-object grasping addresses how to grasp multiple objects simultaneously, typically utilizing all the DoF of the hand [13].
We formulate the sequential multi-object grasping problem as generating a sequence of $N$ grasps $\mathcal{G}=\{\mathbf{g}_{i}\}_{i=1}^{N}$ for picking a sequence of $N$ objects $\mathcal{O}=\{\mathbf{O}_{i}\}_{i=1}^{N}$, where each $\mathbf{g}_{n}\in\mathcal{G}$ is restricted to a specific subset $\mathcal{OS}_{n}$ of the hand’s total DoF and $N\geq 2$}. Mathematically, this can be described as
@@
where $\mathcal{G}_{n-1}=\{\mathbf{g}_{i}\}_{i=1}^{n-1}$, $\mathcal{O}_{n-1}=\{\mathbf{O}_{i}\}_{i=1}^{n-1}$, $\mathcal{G}_{0}=\varnothing$, and $\mathcal{O}_{0}=\varnothing$. $E$ in Eq. 1 is a differentiable function that quantifies how well grasp $\mathbf{g}_{n}$ can pick object $\mathbf{O}_{n}$ with the DoF $\mathcal{OS}_{n}$ given all previously generated grasps $\mathcal{G}_{n-1}$ and objects $\mathcal{O}_{n-1}$. We assume all previously grasped objects $\mathcal{O}_{n-1}$ remain fixed in the hand when generating a grasp on the next object $\mathbf{O}_{n}$.
In this work, we represent $\mathcal{OS}_{n}$ as an opposition space (Section IV-A), each object $\mathbf{O}\in\mathcal{O}$ as a triangular mesh, and each grasp $\mathbf{g}\in\mathcal{G}$ as a vector $\mathbf{g}=\left[\mathbf{p},\mathbf{r},\bm{\theta}\right]\in\mathbb{R}^{9+K}$, where $\mathbf{p}\in\mathbb{R}^{3}$ is the hand’s base position, $\mathbf{r}\in\mathbb{R}^{6}$ is the hand’s base orientation in a 6D continuous reprsentation [29], and $\bm{\theta}\in\mathbb{R}^{K}$} is the $K$-dimensional hand joint angles which are 16 for the Allegro Hand. We assume the shape of all objects in $\mathcal{O}$ to be fully known. Next, we will introduce SeqGrasp our algorithm for solving Eq. 1.
@@ IV-A Opposition Space Selection Strategy
The primary objective in sequential multi-object grasping is to maximize the hand’s remaining DoF after each grasp. For this purpose, we propose a grasp planning strategy guided by OSes [30, 11, 31, 25]. An OS is a functional subspace within the hand’s kinematic structure formed by pairs of opposing surfaces (such as fingertips, lateral surfaces of fingers, or palm surfaces) along with the joints that control these surfaces [25]. It represents regions where opposing forces can be applied to create stable grasps. The number of OSes are hand-dependent and vary based on the kinematic structure. Fig. 3(a) shows the seven different OSes for the Allegro Hand.
@@ IV-B Optimization-based Grasp Generation
The next step in the algorithm (Lines 1-1) is to generate a stable, physically plausible, and collision-free grasp that respects the sampled $\mathcal{OS}_{n}$. To achieve this, we formulate $E$ in Eq. 1 as an energy function (Section IV-C) and numerically optimize it using the Metropolis-Adjusted Langevin Algorithm (MALA) [32] (Line 1).
In robotic grasping, MALA has been used to optimize single object grasps $\mathbf{g}_{n}$ by iteratively refining  $\mathbf{p}_{n}$, $\mathbf{r}_{n}$ and $\bm{\theta}_{n}$ according to Langevin dynamics [18, 7]. However, we must adapt MALA to sequential multi-object grasping. To this end, we propose a new grasp as $\hat{\mathbf{g}}_{n}\leftarrow\mathbf{g}_{n}-\gamma\left[\mathbf{1},\mathbf{J}_{n}\right]\odot\Delta$, where $\gamma$ is the step size, $\mathbf{1}\in\mathbb{R}^{9}$ is a padding vector to align the length of $\mathbf{J}$ with $\mathbf{g}$, $\Delta=\partial E/\partial\mathbf{g}_{n}$ is the energy gradient, and $\odot$ is the element-wise (Hadamard) product. $\hat{\mathbf{g}}_{n}$ is accepted if $\alpha\geq u$, where $u\sim\mathcal{U}([0,1])$} and
@@
The above procedure is repeated for a fixed number of steps where, at each step, $\{\mathbf{x}_{j}\}_{j=1}^{2}$ is re-sampled with probability $p_{\text{accept}}$ (Line 1). This resampling process helps accelerate convergence and escape from local minimas [18, 7].
@@ IV-C Energy Function
The force-closure term ($E_{\text{fc}}$) encourages the grasp to be in force-closure equilibrium [33]. The definition for ($E_{\text{fc}}$) originates from [7], where zero friction and uniform contact force magnitudes are assumed because it significantly speeds up the computation. We follow the same definition and write it as
@@ V Dataset Generation
We use SeqGrasp to generate our large-scale dataset SeqDataset  [^2] The dataset is available at https://yulihn.github.io/SeqGrasp/. containing 4.9 million sequential Allegro Hand grasps on over 509 objects from DexGraspNet [7]. Each object is resized to fit within the Allegro Hand by scaling its axis-aligned bounding box to be between 0.06 and 0.10 meters. Then, from the 509 resized objects, we generate 600 unique object sets containing four objects each. The objects in these sets are randomly permuted four times, resulting in 2,400 unique object sequences.
We run Algorithm 1 on the 2,400 unique object sequences with $\mathbf{w}=[50,50,w_{\text{hop}},5,1,5]^{T}$, where $w_{\text{hop}}$, that penalizes hand-object penetration, grows linearly from $5$ to $5e^{2}$. The optimization runs for 6,000 iterations per grasp. We validate the generated grasp sequences in the physics simulator Isaac Gym following the setup in [7]. This setup initializes all object densities to 500 $\text{kg/m}^{3}$, the friction coefficient to 2.0, the hand to the generated grasp configuration, and objects as free-floating. Then, the hand closes until contact is established with the object, which happens when the distance between the hand and the object is less than 2 mm. Finally, a grasp stability test is evaluated by linearly applying an acceleration of 9.8 $\text{m}/\text{s}^{2}$ in all six orthogonal directions for 100 consecutive simulation steps. A grasp sequence is successful if all grasped objects remain in contact with the hand after the grasp stability test and the maximum penetration depth is less than 1 cm. Otherwise, the grasp sequence is a failure. This procedure ultimately produced 4.9 million grasps, of which 870K (17.82%) were successful. Only successful grasps were retained in SeqDataset. The entire data generation process requires approximately 2,500 GPU hours on an NVIDIA A100.
The statistics of SeqDataset are presented in Table II. The results demonstrate that grasps using palm opposition achieved significantly higher success rates, suggesting that these OSes are important in sequential multi-object grasping. Notably, grasps using thumb-index and thumb-palm OSes display lower success rates than other OSes, which deviates from previous research findings that underscore the thumb’s dominant role in human hand manipulation tasks [34]. We hypothesize that this discrepancy may stem from the biomechanical differences between the Allegro and the human hand. Finally, SeqDataset’s grasp consumption, which indicates no more available OSes ($\mathcal{OS}=\emptyset$), aligns with its objective of supporting multi-object grasping, with a significant portion of cases (96.12%) showing potential for sequential grasping of three to four objects, indicating efficient utilization of the hand’s DoF.

### v5

@@ I Introduction
Dexterous grasping has been extensively studied in robotics [bohg2013data, newbury2023deep]. The field has evolved through two primary research directions. First, numerous methods have been developed for generating robotic grasps [dexdiffuser, lu2023ugg, jensdgcc, wei2024dro, dexgraspnet, li2022gendexgrasp], from analytical approaches [dexgraspnet] to data-driven generative models including Generative Adversarial Networks [jensdgcc], diffusion models [lu2023ugg, dexdiffuser], and Variational Autoencoders [li2022gendexgrasp, wei2024dro]. Second, a complementary line of research has focused on understanding human grasping behaviors, including intention recognition [song2013predicting], real-time tracking of hand-object interactions [romero2010hands], grasp taxonomy classification [feix2015grasp], and learning task-oriented grasping strategies from human demonstrations [kokic2020learning].
The primary strategy in dexterous grasp generation has been to bring the robotic hand close to the object and then simultaneously envelop it with all fingers. While this strategy often results in efficient and successful grasp generation, it simplifies dexterous grasping to resemble parallel-jaw grasping, thereby underutilizing the many DoF of multi-fingered robotic hands [dexgraspnet]. In contrast, grasping multiple objects with a robotic hand, particularly in a sequential manner that mirrors human-like dexterity, as shown in Fig. 1, is still an unsolved problem.
@@
We experimentally evaluate SeqGrasp, SeqDiffuser, and the state-of-the-art simultaneous multi-object grasping method MultiGrasp [multigrasp] in simulation and on physical hardware. The simulation results revealed that SeqGrasp and SeqDiffuser perform on par with MultiGrasp for picking one or two objects while outperforming it when picking three to four objects. Moreover, SeqDiffuser demonstrates superior efficiency, generating 256 grasps within one second compared to approximately 1000 seconds for SeqGrasp and MultiGrasp. For the real-world experiments, we replicated the grasp sequences proposed by the methods on a real Allegro Hand attached to a Franka Panda Robot. These results align with the simulation findings, demonstrating that SeqGrasp reaches a 43.33% higher grasp success rate than MultiGrasp.
@@
MOG [sun2022multi] Human&Barrett Sim./Real 11 \ 28K Simu./Seq. \
Grasp’Em [multigrasp] Shadow Sim. 8 36 90K Simultaneous 2
@@ II-A Dexterous Grasping
Analytical methods. Early research in dexterous grasping generated stable grasps by optimizing a grasp quality metric such as the force-closure metric [graspit, old3]. Although these methods are theoretically sound, they are computationally demanding because (i) the many DoF dexterous hands cause high-dimensional search spaces [eigengrasp], and (ii) the quality metrics are expensive to compute [dfc]. Consequently, some methods have focused on reducing the search space by imposing constraints on the hand [old1, old2] or restricting it to joint configuration subspaces [eigengrasp]. Another line of work has proposed a computationally cheap differentiable force-closure estimator [dfc, dexgraspnet], which has the advantage of being hand-agnostic. This work extends the differentiable force-closure measure from [dexgraspnet] to sequential multi-object grasping.
Data-driven methods. Recent advancements in machine learning have significantly improved dexterous grasp generation [dexdiffuser, mayer2022ffhnet, lu2023ugg, wei2024dro]. Nowadays, deep generative models can generate thousands of dexterous grasps on previously unseen and partially observed objects within seconds [dexdiffuser, mayer2022ffhnet, lu2023ugg, wei2024dro], something analytical methods cannot. Still, only a few data-driven dexterous grasping methods have been developed for multi-object grasping [he2025sequential, multigrasp]. In this work, we train a new diffusion-based sequential multi-object grasp sampler SeqDiffuser inspired from [dexdiffuser] on our own optimization-generated sequential multi-object dataset SeqDataset.
@@ II-B Multi-Object Grasping
Multi-object grasping presents unique challenges due to the complex multi-object interactions and the high-dimensional configuration space spanned by the hand and the objects. Some prior parallel-jaw multi-object grasping methods [agboh2022multi, yonemaru2025learning] explored multi-object push grasps where scattered objects are first pushed together to facilitate multi-object grasping. However, these methods are limited by their reliance on simple shape primitives and parallel-jaw grippers. In comparison, our work can handle objects of diverse shapes and sizes.
A few works have addressed dexterous multi-object grasping [multigrasp, he2025sequential, yao2023exploiting] where [multigrasp] targets simultaneous multi-object grasping while [he2025sequential, yao2023exploiting] targets sequential multi-object grasping. Li et al. [multigrasp] proposed MultiGrasp a two-stage simultaneous multi-object dexterous grasping framework where a generative grasp sampler proposed poses to simultaneously pick many objects, followed by a learned policy for executing the pick. Simultaneously grasping multiple objects is an efficient strategy because it avoids repeated hand repositioning, which reduces total execution time, but the main limitation is that objects must be spatially close and of similar size and shape. In comparison, our method can handle scattered objects of different shapes and sizes by sequentially picking one at a time. The other works that do sequential multi-object grasping [he2025sequential, yao2023exploiting] restrict the grasping to a maximum of two objects [he2025sequential] or to primitive object shapes such as cylinders or spheres [yao2023exploiting]. In comparison, our method can handle up to four objects of complex shapes and sizes. We also use our method to collect the largest sequential multi-object grasping dataset to date.
@@ II-C Datasets for Dexterous Grasping
Large-scale dexterous grasp datasets [jensdgcc, dexgraspnet, chen2024bodex, sun2022multi, multigrasp, mayer2022ffhnet, graspxl, realdex] have significantly advanced the training of data-driven dexterous grasping methods. However, most of these datasets target single-object grasping [jensdgcc, dexgraspnet, chen2024bodex, mayer2022ffhnet, graspxl, realdex], with only a few for multi-object grasping [sun2022multi, multigrasp]. As shown in Table I, these existing multi-object datasets are small and predominantly focus on simultaneous rather than sequential grasping. Therefore, we collect the new large-scale sequential multi-object grasping dataset SeqDataset [^1] The dataset is available at https://yulihn.github.io/SeqGrasp/. using our method SeqGrasp. SeqDataset is, to date, by far the largest multi-object grasping dataset.
@@ III Problem Formulation
To contrast, simultaneous multi-object grasping addresses how to grasp multiple objects simultaneously, typically utilizing all the DoF of the hand [multigrasp].
We formulate the sequential multi-object grasping problem as generating a sequence of $N$ grasps $\mathcal{G}=(\mathbf{g}_{i})_{i=1}^{N}$ for picking a sequence of $N$ objects $\mathcal{O}=(\mathbf{O}_{i})_{i=1}^{N}$, where $N\geq 2$} and $(\cdot)$ denotes an ordered sequence. Each $\mathbf{g}_{n}\in\mathcal{G}$ is restricted to a specific subset $\mathcal{OS}_{n}$ of the hand’s total DoF. Mathematically, this can be described as
@@
where $\mathcal{G}_{n-1}=(\mathbf{g}_{i})_{i=1}^{n-1}$, $\mathcal{O}_{n-1}=(\mathbf{O}_{i})_{i=1}^{n-1}$, $\mathcal{G}_{0}=\varnothing$, and $\mathcal{O}_{0}=\varnothing$. $E$ in Eq. 1 is a differentiable function that quantifies how well grasp $\mathbf{g}_{n}$ can pick object $\mathbf{O}_{n}$ with the DoF $\mathcal{OS}_{n}$ given all previously generated grasps $\mathcal{G}_{n-1}$ and objects $\mathcal{O}_{n-1}$. We assume all previously grasped objects $\mathcal{O}_{n-1}$ remain fixed in the hand when generating a grasp on the next object $\mathbf{O}_{n}$.
In this work, we represent $\mathcal{OS}_{n}$ as an opposition space (Section IV-A), each object $\mathbf{O}\in\mathcal{O}$ as a triangular mesh, and each grasp $\mathbf{g}\in\mathcal{G}$ as a vector $\mathbf{g}=\left[\mathbf{p},\mathbf{r},\bm{\theta}\right]\in\mathbb{R}^{9+K}$, where $\mathbf{p}\in\mathbb{R}^{3}$ is the hand’s base position, $\mathbf{r}\in\mathbb{R}^{6}$ is the hand’s base orientation in a 6D continuous reprsentation [6drot], and $\bm{\theta}\in\mathbb{R}^{K}$} is the $K$-dimensional hand joint angles which are 16 for the Allegro Hand. We assume the shape of all objects in $\mathcal{O}$ to be fully known. Next, we will introduce SeqGrasp our algorithm for solving Eq. 1.
@@ IV-A Opposition Space Selection Strategy
The primary objective in sequential multi-object grasping is to maximize the hand’s remaining DoF after each grasp. For this purpose, we propose a grasp planning strategy guided by OSes [iberall1986opposition, feix2015grasp, smeets2019review, yao2023exploiting]. An OS is a functional subspace within the hand’s kinematic structure formed by pairs of opposing surfaces (such as fingertips, lateral surfaces of fingers, or palm surfaces) along with the joints that control these surfaces [yao2023exploiting]. It represents regions where opposing forces can be applied to create stable grasps. The number of OSes are hand-dependent and vary based on the kinematic structure. Fig. 3(a) shows the seven different OSes for the Allegro Hand.
@@ IV-B Optimization-based Grasp Generation
The next step in the algorithm (Lines 1-1) is to generate a stable, physically plausible, and collision-free grasp that respects the sampled $\mathcal{OS}_{n}$. To achieve this, we formulate $E$ in Eq. 1 as an energy function (Section IV-C) and numerically optimize it using the Metropolis-Adjusted Langevin Algorithm (MALA) [Roberts1996ExponentialCO] (Line 1).
In robotic grasping, MALA has been used to optimize single object grasps $\mathbf{g}_{n}$ by iteratively refining  $\mathbf{p}_{n}$, $\mathbf{r}_{n}$ and $\bm{\theta}_{n}$ according to Langevin dynamics [dfc, dexgraspnet]. However, we must adapt MALA to sequential multi-object grasping. To this end, we propose a new grasp as $\hat{\mathbf{g}}_{n}\leftarrow\mathbf{g}_{n}-\gamma\left[\mathbf{1},\mathbf{J}_{n}\right]\odot\Delta$, where $\gamma$ is the step size, $\mathbf{1}\in\mathbb{R}^{9}$ is a padding vector to align the length of $\mathbf{J}$ with $\mathbf{g}$, $\Delta=\partial E/\partial\mathbf{g}_{n}$ is the energy gradient, and $\odot$ is the element-wise (Hadamard) product. $\hat{\mathbf{g}}_{n}$ is accepted if $\alpha\geq u$, where $u\sim\mathcal{U}([0,1])$} and
@@
The above procedure is repeated for a fixed number of steps where, at each step, $\{\mathbf{x}_{j}\}_{j=1}^{2}$ is re-sampled with probability $p_{\text{accept}}$ (Line 1). This resampling process helps accelerate convergence and escape from local minimas [dfc, dexgraspnet].
@@ IV-C Energy Function
The force-closure term ($E_{\text{fc}}$) encourages the grasp to be in force-closure equilibrium [rimon1996force]. The definition for ($E_{\text{fc}}$) originates from [dexgraspnet], where zero friction and uniform contact force magnitudes are assumed because it significantly speeds up the computation. We follow the same definition and write it as
@@ V Dataset Generation
We use SeqGrasp to generate our large-scale dataset SeqDataset  [^2] The dataset is available at https://yulihn.github.io/SeqGrasp/. containing 4.9 million sequential Allegro Hand grasps on over 509 objects from DexGraspNet [dexgraspnet]. Each object is resized to fit within the Allegro Hand by scaling its axis-aligned bounding box to be between 0.06 and 0.10 meters. Then, from the 509 resized objects, we generate 600 unique object sets containing four objects each. The objects in these sets are randomly permuted four times, resulting in 2,400 unique object sequences.
We run Algorithm 1 on the 2,400 unique object sequences with $\mathbf{w}=[50,50,w_{\text{hop}},5,1,5]^{T}$, where $w_{\text{hop}}$, that penalizes hand-object penetration, grows linearly from $5$ to $5e^{2}$. The optimization runs for 6,000 iterations per grasp. We validate the generated grasp sequences in the physics simulator Isaac Gym following the setup in [dexgraspnet]. This setup initializes all object densities to 500 $\text{kg/m}^{3}$, the friction coefficient to 2.0, the hand to the generated grasp configuration, and objects as free-floating. Then, the hand closes until contact is established with the object, which happens when the distance between the hand and the object is less than 2 mm. Finally, a grasp stability test is evaluated by linearly applying an acceleration of 9.8 $\text{m}/\text{s}^{2}$ in all six orthogonal directions for 100 consecutive simulation steps. A grasp sequence is successful if all grasped objects remain in contact with the hand after the grasp stability test and the maximum penetration depth is less than 1 cm. Otherwise, the grasp sequence is a failure. This procedure ultimately produced 4.9 million grasps, of which 870K (17.82%) were successful. Only successful grasps were retained in SeqDataset. The entire data generation process requires approximately 2,500 GPU hours on an NVIDIA A100.
The statistics of SeqDataset are presented in Table II. The results demonstrate that grasps using palm opposition achieved significantly higher success rates, suggesting that these OSes are important in sequential multi-object grasping. Notably, grasps using thumb-index and thumb-palm OSes display lower success rates than other OSes, which deviates from previous research findings that underscore the thumb’s dominant role in human hand manipulation tasks [thumb]. We hypothesize that this discrepancy may stem from the biomechanical differences between the Allegro and the human hand. Finally, SeqDataset’s grasp consumption, which indicates no more available OSes ($\mathcal{OS}=\emptyset$), aligns with its objective of supporting multi-object grasping, with a significant portion of cases (96.12%) showing potential for sequential grasping of three to four objects, indicating efficient utilization of the hand’s DoF.